OpenStudy (anonymous):

Evaluate the expression (2x^3y^2)^2 for x=2 and y=3 a. 10,368 b. 20,736 c. 23,328 d. 46,656

OpenStudy (anonymous):

\((2x^3y^2)^2\\=(2(2)^3(3)^2)^2\\=?\) which part are you having difficulty with?

OpenStudy (anonymous):